MULTIPLE REGRESSION

*****PLEASE HELP ME WITH PART Eresearcher emplo e turinde ender variables in a re ession mnde to reti a measure cf hie navel y eve a vaca n ns c osen by 4 8 AT THE BOTTOM*****

PLEASE HELP ME UNDERSTAND HOW SOLVE PROBLEM IF YOU KNOW HOW TO DO MULTIPLE REGRESSION
Answer #1

Solution:
Given:

N = Total sample size = 408

k = number of independent variables = 4

Level of significance = \alpha = 0.025

Part e) Find rejection region

We have to find F critical value.

We use following excel command:

=F.DIST.RT( probability , df_numerator , df_denominator)

Here

df_numerator= k = 4

df_denominator = N - k - 1 = 408 - 4 - 1 = 403

Thus

=F.DIST.RT( 0.025 , 4 , 403 )

=2.82

Thus F critical value = 2.82

Thus rejection region is:

F > 2.82